OpenStudy (anonymous):

Give the excluded value(s) for the following fraction. y+5/y^2+4y-32

OpenStudy (anonymous):

you can't divide by 0, which is means that denominator can not be =0 solve: y^2+4y-32=0 find y - those values are excluded can you finish it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

y^2+4y-32=(y+8)*(y-4)=0 how about now?

OpenStudy (anonymous):

i havent even learned how to do this

OpenStudy (anonymous):

if a*b=0 then a=0 or b=0... right? so from the above: y+8=0 or y-4=0 so y=...?

OpenStudy (anonymous):

8,-4

OpenStudy (anonymous):

y=-8 and y=4 are excluded (you forgot to change the sign...)